OpenStudy (anonymous):

Find a such that f(x)=ax^2+5x+6 has a minimum value of 7(1/24)

OpenStudy (anonymous):

vertex is at x = -5/(2a), so f(-5/(2a)) = 7/24 = a(-5/(2a))^2+5(-5/(2a))+6 solve for a, and a = 150/137

OpenStudy (anonymous):

thanks :)